OpenStudy (anonymous):

Work out the value of 4 ½ - 2 ⅙ Give your answer as a mixed number in its simplest form and show your working out

OpenStudy (anonymous):

(4*2+1)/2 =9/2 (2*6+1)/6 = 13/6 9/2 - 13/6 = (27-13)/6 = 14/6 = 7/2

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

emunrradtvamg was almost correct. When you divide 14/6, (divide both sides by 2) you should get 7/3 not 7/2 7/3 as a mixed number is 2 1/3